OpenStudy (anonymous):

A set of equations is given below: Equation S: y = x + 9 Equation T: y = 2x + 1 Which of the following steps can be used to find the solution to the set of equations? x = 2x + 1 x + 9 = 2x x + 1 = 2x + 9 x + 9 = 2x + 1

OpenStudy (lynfran):

equate them

OpenStudy (anonymous):

I know the answer but I can't explain. Sorry

OpenStudy (lynfran):

example y=3x+1 y=5x+2 since both have y=to... we equate them 3x+1=5x+2

OpenStudy (lynfran):

by doing that we can solve for the variable "x"

OpenStudy (anonymous):

by this you mean x + 9 = 2x + 1? I was thinking its the last one because we have to you know equate them but i dont know how

OpenStudy (lynfran):

correct

OpenStudy (anonymous):

Ok thankyou for helping me

OpenStudy (lynfran):

welcome